OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (−3, −1) with a slope of 4. y = 4x − 11 y = 4x + 11 y = 4x + 7 y = 4x – 7 Which of the following statements best describes the effect of replacing the graph of f(x) with the graph of f(x) − 5? The graph shifts 5 units up. The graph shifts 5 units down. The graph shifts 5 units left. The graph shifts 5 units right.

OpenStudy (anonymous):

You doing FLVS?

OpenStudy (anonymous):

yep homeschooling

OpenStudy (anonymous):

What lesson?

OpenStudy (anonymous):

Math Module 3 3.09

OpenStudy (anonymous):

can someone viewing this question help me answer these two questions plz

OpenStudy (anonymous):

Ok for first question, find out where (-3, -1) can fit in those equations.

OpenStudy (anonymous):

???

OpenStudy (anonymous):

Just plug in -3 for x and -1 for y, and see if they match

OpenStudy (anonymous):

I got -12 +b = -1

OpenStudy (anonymous):

Here, let's do the first one: y = 4x − 11 -1 = 4(-3) - 11 -1 = -12 -11 -1 does not equal to -23 So A cannot be the answer

OpenStudy (anonymous):

where did the 11 come from?

OpenStudy (anonymous):

11 was there in that equation from the options you gave, remember?

OpenStudy (anonymous):

oh I see

OpenStudy (anonymous):

Yeah, so A cannot be the answer. Try doing B now

OpenStudy (anonymous):

B is correct

OpenStudy (anonymous):

I got -1 = -1

OpenStudy (anonymous):

Yup :)

OpenStudy (anonymous):

thx :)

OpenStudy (anonymous):

No problem. Now let's do the second question

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I think the answer is that it shifts down

OpenStudy (anonymous):

Yeah I think so too

OpenStudy (anonymous):

So answer should be B

OpenStudy (anonymous):

ok is their proof to see if this is right. I need to get a 70% or higher in order to do my DBA with my teacher

OpenStudy (anonymous):

btw I fanned you

OpenStudy (anonymous):

Here check this out https://www.desmos.com/calculator/q9hgto7zdn

OpenStudy (anonymous):

k thanks. I also want to ask if their is a straight line on the graph is the slope undefined or 0?

OpenStudy (anonymous):

The slope is undefined when it is a vertical line. 0 is when it is horizontal

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

kk see you around and thanks so much. I'll let you know if I got them right :D